# SynCL: A Synergistic Training Strategy with Instance-Aware Contrastive Learning for End-to-End Multi-Camera 3D Tracking

### v2

@@ 3.5 Instance-aware Contrastive Learning
In the tracking-by-attention paradigm, track queries, serving as feature priors, need to associate and update persistent targets within the current frame. However, object queries struggle to effectively transfer the knowledge of persistent targets learned by detection pre-training to track queries due to self-attention mechanisms. In contrast, our proposed parallel decoder employs the hybrid label assignment, enabling object queries and track queries to share identical ground-truth matching. Consequently, track queries naturally acquire high-quality learning samples. We treat object and track queries assigned to the same instance as positive samples and the remaining query pairs as negative samples. We utilize kernel-based contrastive learning to alleviate the problem of insufficient pairs Tsai et al. (2022), which is defined as:
$\mathcal{L}_{CL}=-\log\frac{\exp(\mathcal{K}[e^{obj}_{i},e^{trk}_{j}])}{\sum_{k=1,k\neq j}^{N}\exp(\mathcal{K}[e^{obj}_{i},e^{trk}_{k}])}$ (11)
where $e^{obj}$ and $e^{trk}$ are the embedding vectors of object and track queries in the C-decoder. $\mathcal{K}[\cdot]$ refers to the kernel function mapping with $\phi(\cdot)$ and $g_{\theta}(\cdot)$ shown as:
@@
where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ is the Hadamard product. In practice, we employ $\phi(\cdot)$ and $g_{\theta_{obj}}(\cdot)$ as a multi-layer perceptron (MLP) and a FFN. We do not utilize any projection for track queries for simplicity. The overall process can be written as:
@@
Instance-aware contrastive learning aims to enable the learning of enriched representations from both types of queries in a bidirectional manner, effectively bridging the gap between detection and tracking.
@@ 4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics
Method Detector Tracking Detection
@@ 4.1.1 Datasets and Evaluation Metrics
MUTR3D Ding et al. (2024) DETR3D 32.1% 1.448 45.2% 28.3% 474 15269 43828 - -
SynCL(ours) DETR3D 35.8% 1.391 49.2% 32.9% 588 14311 40740 - -
PF-Track-S Pang et al. (2023) PETR 40.8% 1.343 50.7% 37.6% 166 15288 40398 47.7% 37.8%
SynCL (ours) PETR 44.7% 1.262 56.5% 40.8% 203 15344 36801 49.7% 39.6%
TBA-Baseline PETRv2 43.2% 1.272 55.0% 40.6% 173 14106 37065 50.4% 41.0%
SynCL (ours) PETRv2 45.7% 1.260 56.8% 43.0% 170 13411 36756 51.1% 42.0%
@@
SynCL (ours) PETR 48.5% 1.218 58.7% 44.5% 223 13518 34107 51.3% 42.7%
CC-3DT Fischer et al. (2022) BEVFormer 41.0% 1.274 53.8% 35.7% 3334 - - - -
DQTrack Li et al. (2023) PETRv2 44.6% 1.251 - - 1193 - - 50.3% 41.0%
ADA-Track Ding et al. (2024) PETR 47.9% 1.246 60.2% 43.0% 767 15385 31402 - -
PF-Track-F Pang et al. (2023) PETR 47.9% 1.227 59.0% 43.5% 181 16149 32778 51.4% 42.2%
Table 1: Comparison results on nuScenes val set. For a fair comparison, we divide the experiment into four groups based on the baseline settings. Our method is shown in light gray background. $\dagger$ indicates we omit the Past reasoning module in PF-Track.
We evaluate the proposed SynCL on nuScenes Caesar et al. (2020), a large-scale benchmark for autonomous driving, that contains 700, 150, and 150 scenes for training, validation, and testing, respectively. The dataset covers 10 types of common objects on the road. For the tracking task, nuScenes selects a subset of 7 movable categories, excluding static objects such as traffic cones. Consistent with official evaluation metrics for 3D MOT, we primarily report AMOTA (average multi-object tracking accuracy) and AMOTP (average multi-object tracking precision) Weng et al. (2020b). We also report metrics of NDS (nuScenes detection score) and mAP (mean Average Precision) for detection to fully evaluate multi-task performance.
4.1.2 Baseline and Implementation Details
To assess the generality of our proposed synergistic training strategy, we integrate SynCL within several MOT works using different detectors. Specifically, our experiments involve three query-based detectors: DETR3D Wang et al. (2021), PETR Liu et al. (2022), and PETRv2 Liu et al. (2023). For DETR3D and PETR, we follow the experimental setups in ADA-Track Ding et al. (2024) and PF-Track Pang et al. (2023), respectively. Notably, PF-Track includes both a full resolution ($1600\times 640$, indicated with ”-F”) and a small resolution ($800\times 320$, indicated with ”-S”) setting. Due to GPU memory constraints, we reproduce PF-Track-F but omit the Past reasoning module. Regarding PETRv2, we establish the corresponding TBA-Baseline by substituting only the detector in the PF-Track-S framework. For the ablation studies, configurations of PF-Track-S are employed as the default. All methods are trained using the AdamW Loshchilov and Hutter (2019) optimizer with a weight decay of $1.0\times 10^{-2}$. The learning rate is initially set to $2.0\times 10^{-4}$ and following a cosine annealing schedule.
id o2m o2o CL Tracking Detection
@@ 4.1.2 Baseline and Implementation Details
1 40.8% 1.343 47.7% 37.8%
2 42.2% 1.327 49.8% 38.9%
3 40.9% 1.339 48.1% 37.9%
4 42.6% 1.286 49.9% 39.9%
5 44.7% 1.262 49.7% 39.6%
Table 2: Component effectiveness of SynCL. ”o2m”, ”o2o” and ”CL” denote the one-to-many, one-to-one and contrastive learning for object and track queries in C-decoder.
4.2 Main results
We evaluate SynCL on the nuScenes validation set. Tab.1 presents the comparative results after integrating SynCL into various trackers, with each baseline shown at the top of each detector group. The experimental results lead to four key observations: 1) SynCL can be adapted to various tracking-by-attention paradigm methods, including MUTR3D and PF-Track. Specifically, SynCL improves MUTR3D, PF-Track-S, PF-Track-F by $+3.7\%$, $+3.9\%$, and $+3.5\%$ AMOTA, respectively. 2) SynCL exhibits compatibility with various detectors utilizing different attention variants, indicating that the issues associated with the self-attention mechanism are widespread. Building upon our TBA-Baseline, SynCL achieves additional $+2.5\%$ boosts in AMOTA, outperforming the decouple method DQTrack by 1.1% AMOTA. 3) As a synergistic training strategy, SynCL uniquely enhances both detection and tracking performance, providing a novel solution for end-to-end tracking. Notably, SynCL delivers 2% NDS improvements compared to the PF-Track-S.
Compared to the current state-of-the-art works, SynCL with PETR at the small resolution surpasses CC-3DT Fischer et al. (2022) at the full resolution using a stronger BEVFormer detector by $+3.7\%$ AMOTA. With lower inference latency and model complexity, SynCL achieves 48.5% AMOTA, outperforming both PF-Track and ADA-Track.
4.3 Ablation Studies
4.3.1 Component effectiveness
@@
In Tab. 2, we analyze the importance of each proposed component in SynCL for the model performance. According to the source of supervision, we decompose SynCL into one-to-many learning for object queries, one-to-one learning for track queries and contrastive learning for both queries in the C-decoder. In general, our proposed components enjoy consistent performance improvements. Specifically, by introducing hybrid matching it outperforms the baseline by 1.8% AMOTA. Further integrating the instance-aware contrastive learning brings an extra 2.1% improvement in AMOTA. Together, our proposed components complement each other to effectively address multi-task learning challenges.
@@
ATSS 39.3% 1.367 49.9% 47.0% 37.3%
SimOTA 42.8% 1.323 52.5% 49.2% 38.8%
DETA 43.1% 1.320 54.2% 49.4% 38.6%
Ours 44.7% 1.262 56.5% 49.7% 39.6%
Table 3: Comparison results of different one-to-many label assignment methods in C-decoder.
Query source Matching Range Strategy Tracking Detection
@@
S-decoder New-born GTs one-to-many 38.8% 1.357 52.4% 15461 39930 47.7% 38.8%
@@
Table 4: Comparision results of different label assignment variants for SynCL.
Figure 5: Comparative analysis of training cost compared to baseline with training time (h) and GPU memory (G).
4.3.2 Computation and memory complexity
We implement the parallel decoder in an efficient way following Chen et al. (2023). As a result, SynCL only takes an acceptable increase in training cost including GPU memory and training time as shown in Fig.5. For example, with TBA-Baseline and MUTR3D, training time is increased by 0.3 and 0.4 hours per epoch. The memory increases are just 4G and 1G, respectively. Thus, the weight-shared decoder design maintains the total number of parameters and slightly increases the computational cache.
@@ 4.3.2 Computation and memory complexity
Table 5: Ablation of maximum assignment constraint $K$.
4.3.3 Variants of query filtering
We examine the impact of different one-to-many assignment methods compared with our proposed GMM-based filtering module in Tab. 3. Concretely, ATSS Zhang et al. (2020), SimOTA Ge et al. (2021), DETA Zhang et al. (2022a) are chosen as the alternatives. Both ATSS and SimOTA dynamically compute the assignment count for each ground truth based on the IoU matrix. However, ATSS proves overly restrictive for 3D perception tasks, consequently compromising performance. The fixed-threshold method of DETA proves insufficient for mining potential high-quality candidates. We demonstrate the candidate selection effect of our GMM-based filtering module in Fig. 6.
Figure 6: An illustration of Cost-based Query Filtering.
4.3.4 Variants of label assignment
In Tab. 4, we analyze the impact of different label assignment strategies for object queries on tracking and detection performance. We consider three different variants from the perspectives of query source, matching range, and matching strategy, compared to our baseline (highlighted in gray). Based on the experimental results we draw the following three conclusions: 1) The one-to-many assignment is not applicable to the standard decoder due to the NMS effect of the self-attention mechanism. 2) Object queries in the C-decoder cannot benefit from a broader matching range with all ground-truth targets. The reduction of positive samples proportion in contrastive learning may diminish the effectiveness of synergistic learning. 3) One-to-one assignment exhibits suboptimal alignment with the rich prediction properties of only cross-attention in C-decoder, leading to limited improvements.
4.3.5 Hyperparameters
We study two types of hyperparameters in our approach, the maximum assignment constraint $K$ per GT in Eq.9 and the training sample length $T$. For $K$, according to Tab.5, the best performance is achieved when $K=5$. Possibly, a smaller matching number could lack supervision signals, while a larger matching number could introduce noisy and unreliable candidates. For $T$ in Tab.6, we observe that when the frame number is increased from 2 to 3, SynCL obtains a boost of $+1.6\%$ AMOTA. However, when $T>3$, SynCL cannot improve the model performance further. We infer that more consecutive frames lead to more optimization objectives, making the training gradient increase, thus affecting the training stability. Considering the performance and computational cost, we use $K=5$ and $T=3$ as the default setting.
@@ 4.3.5 Hyperparameters
Table 6: Ablation of training sample length $T$.

### v3

@@ 3.5 Instance-aware Contrastive Learning
To further transfer object queries’ knowledge of persistent targets learned by one-to-many assignment in Hybrid Matching to track queries, we employ contrastive learning for object queries and their corresponding track queries with identical ground-truth matching. As shown in Fig. 4, contrastive learning pulls representations of object queries closer to those of track queries assigned to the same instance, while pushing them further apart from representations of distinct instances in the latent space. Unlike self-supervised learning methods [32, 33], which involve a large number of sample pairs (e.g., 4096), there are typically fewer than 100 track queries propagated from the previous frame to serve as samples pairs. To alleviate this issue of insufficient sample pairs, we draw on [34] and utilize kernel-based contrastive learning, which is formulated as:
$\mathcal{L}_{CL}(\mathcal{K};\tau)=-\log\frac{\exp(\mathcal{K}[e^{obj}_{i},e^{trk}_{j}]/\tau)}{\sum_{k=1,k\neq j}^{N}\exp(\mathcal{K}[e^{obj}_{i},e^{trk}_{k}]/\tau)}$ (11)
where $e^{obj}$ and $e^{trk}$ are the embeddings of object and track queries in the C-decoder. $\tau$ is a temperature hyper-parameter [35]. $\mathcal{K}[\cdot]$ refers to the kernel function mapping with $\phi(\cdot)$ and $g_{\theta}(\cdot)$:
@@
where $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ is the Hadamard product. In practice, we employ $\phi(\cdot)$ and $g_{\theta_{obj}}(\cdot)$ as a multi-layer perceptron (MLP) and a FFN. We do not utilize any projection for track queries for simplicity:
@@
We apply Instance-aware Contrastive Learning only in the training stage. During inference, the additional model parameters used for contrastive learning from Eq. 14 are discarded.
@@ 4.1 Experimental Setup
Datasets and Evaluation Metrics. We evaluate the proposed SynCL on nuScenes [36], a large-scale benchmark for autonomous driving, which contains 700, 150, and 150 scenes for training, validation, and testing, respectively. The dataset covers 10 types of common objects on the road. For the tracking task, nuScenes selects a subset of 7 movable categories, excluding static objects such as traffic cones. Consistent with official evaluation metrics for 3D MOT, we primarily report AMOTA (average multi-object tracking accuracy) and AMOTP (average multi-object tracking precision) [37]. We also report metrics of NDS (nuScenes detection score) and mAP (mean Average Precision) for detection to fully evaluate multi-task performance.
Baseline and Implementation Details. To assess the generality of our proposed synergistic training strategy, we integrate SynCL within PF-Track [5] and MUTR3D [4] frameworks using different detectors. Specifically, our experiments involve four query-based detectors: DETR3D [30], PETR [29], PETRv2 [31] and StreamPETR [26]. For DETR3D and PETR, we follow the experimental setups in ADA-Track [6] and PF-Track, respectively. Regarding PETRv2, we establish the baseline by substituting only the detector in the PF-Track framework referred to as Baseline#1. Unlike Onetrack [8], our experiments with the StreamPETR detector in MUTR3D framework are initialized from pre-trained detector weights and we then train the tracker in the sliding window mode to ensure the temporal gradient flow, denoted as Baseline#2. Following common practice, we set the training epoch to 24 for experiments with resolution of $900/640\times 1600$, while we set it to 12 for experiments with resolution of $320\times 800$. All methods are trained using the AdamW [38] optimizer with a weight decay of $1.0\times 10^{-2}$. The learning rate is initially set to $2.0\times 10^{-4}$ and following a cosine annealing schedule. We conduct all experiments on eight A100 GPUs with a batch size of 1 and each training batch consists of a clip of three consecutive frames from different timestamps. For the ablation studies, configurations of PF-Track with a small resolution $320\times 800$ are employed as the default.
Inference. We claim that SynCL does not employ the weight-shared parallel decoder during the inference stage and does not modify the inference procedures of any baseline, thus having no impact on inference speed. For more details on the runtime analysis, please refer to Appendix B.1.
4.2 Main results
Table 1: Comparison results of adopting SynCL with tracking-by-attention (TBA) baselines on nuScenes val set. We divide the experiment into four groups based on the detector settings. * denotes results from [6]
Method Backbone Detector Resolution Tracking Detection
@@ 4.2 Main results
MUTR3D* [4] R101 DETR3D 900 $\times$ 1600 32.1% 1.448 45.2% 28.3% 474 15269 43828 - -
SynCL (ours) R101 DETR3D 900 $\times$ 1600 35.8% 1.391 49.2% 32.9% 588 14311 40740 - -
PF-Track [5] V2-99 PETR 320 $\times$ 800 40.8% 1.343 50.7% 37.6% 166 15288 40398 47.7% 37.8%
SynCL (ours) V2-99 PETR 320 $\times$ 800 44.7% 1.262 56.5% 40.8% 203 15344 36801 49.7% 39.6%
Baseline#1 [5] V2-99 PETRv2 320 $\times$ 800 43.2% 1.272 55.0% 40.6% 173 14106 37065 50.4% 41.0%
SynCL (ours) V2-99 PETRv2 320 $\times$ 800 45.7% 1.260 56.8% 43.0% 170 13411 36756 51.1% 42.0%
Baseline#2 [4] V2-99 Stream 320 $\times$ 800 49.6% 1.164 57.3% 42.9% 411 13962 33526 57.6% 48.5%
SynCL (ours) V2-99 Stream 320 $\times$ 800 51.8% 1.149 58.8% 45.2% 540 13639 33368 58.7% 49.2%
Generality of SynCL. We first evaluate SynCL on the nuScenes validation set. Tab. 1 presents the comparative results after integrating SynCL into various trackers, with each baseline shown at the top of each detector group. The experimental results lead to four key observations: 1) SynCL can be adapted to various tracking-by-attention paradigm methods. Specifically, SynCL improves MUTR3D and PF-Track by $+3.7\%$ and $+3.9\%$ AMOTA, respectively. 2) SynCL exhibits compatibility with various detectors utilizing different attention variants, indicating that the issues associated with the self-attention mechanism are widespread. Building upon our Baseline#1 and Baseline#2, SynCL achieves additional $+2.5\%$ and $2.2\%$ boosts in AMOTA. 3) SynCL does not significantly increase the number of false positives while consistently improving the rate of recall. This indicates that our dynamic query filtering can select more high-quality candidates. 4) As a synergistic training strategy, SynCL uniquely enhances both detection and tracking performance, providing a novel solution for end-to-end tracking. Notably, SynCL delivers 2% NDS improvements compared to the PF-Track.
State-of-the-art Comparison. In Tab. 2, we first compare SynCL with leading camera-based 3D MOT methods on the nuScenes validation set. With the PETR detector configuration, SynCL achieves the highest performance in both AMOTA and AMOTP. Remarkably, SynCL does not use the additional association module, resulting in lower inference latency compared to the current leader, ADA-Track++ [7]. With the StreamPETR detector configuration, SynCL significantly outperforms the previous state-of-the-art OneTrack by a margin of 4.1% in AMOTA.
Table 2: Comparison with leading camera-based methods on the nuScenes val set. The FPS is measured on a single NVIDIA A100 GPU, from the input images to the final tracking results. * denotes results from [25]
@@
CC-3DT [39] R101 BEVFormer 900 $\times$ 1600 42.9% 1.257 53.4% 38.5% 2219 - - -
DQTrack [11] V2-99 PETRv2 320 $\times$ 800 44.6% 1.251 - - 1193 - - 8.6
MUTR3D* [4] V2-99 PETR 640 $\times$ 1600 44.3% 1.299 55.2% 41.6% 175 11943 36861 6.1
PF-Track [5] V2-99 PETR 640 $\times$ 1600 47.9% 1.227 59.0% 43.5% 181 16149 32778 5.2
ADATrack++ [7] V2-99 PETR 640 $\times$ 1600 50.4% 1.197 60.8% 44.5% 613 14839 30616 3.2
OneTrack [8] V2-99 Stream 640 $\times$ 1600 54.8% 1.088 61.8% 47.9% 389 - - -
SynCL (ours) V2-99 PETR 640 $\times$ 1600 50.7% 1.183 61.3% 46.2% 248 14506 30577 5.2
SynCL (ours) V2-99 Stream 640 $\times$ 1600 58.9% 1.016 64.0% 51.5% 652 13946 27330 5.7
Table 3: Comparison on nuScenes test set. ’E2E’ stands for the end-to-end detection and tracking model.
@@
CC-3DT [39] 41.0% 1.274 53.4% 38.5%
PF-Track [5] 43.4% 1.252 53.8% 37.8%
STAR-Track [24] 43.9% 1.256 56.2% 40.6%
ADATrack++ [7] 50.0% 1.144 59.5% 45.6%
DQTrack [11] 52.3% 1.096 62.2% 44.4%
OneTrack [8] 55.4% 1.021 60.8% 46.1%
DORT [18] 57.6% 0.951 63.4% 48.4%
SynCL (ours) 58.8% 0.976 67.1% 50.4%
Moreover, we present comparison results on nuScenes test set in Tab. 3. SynCL achieves leading performance with 58.8% AMOTA and excels over the state-of-the-art camera-based methods. Notably, SynCL surpasses the dominant two-stage tracker [18] by 1.2% in AMOTA.
Computation and memory complexity. We evaluate the training costs of the models in Tab. 1. We implement the parallel decoder in an efficient way following [40]. As a result, SynCL only takes an acceptable increase in training cost including GPU memory and training time as shown in Fig. 6. For example, with Baseline#1 and MUTR3D, training time is increased by 0.3 and 0.4 hours per epoch. The memory increases are just 4G and 1G, respectively. Thus, the weight-shared decoder design slightly increases the computational cache.
Figure 5: Analysis of training time (h) and GPU memory (G). The inference speed remains unchanged.
Figure 6: Illustration of our Dynamic Query Filtering with Gaussian Mixture Modeling.
4.3 Ablations and Analysis
@@
Component effectiveness. In Tab. 4(a), we analyze the importance of each proposed component in SynCL for the model performance. According to the specific task, we decompose hybrid matching into one-to-many learning for detection and one-to-one learning for tracking. In general, our proposed components enjoy consistent performance improvements. Specifically, by introducing hybrid matching, it outperforms the baseline by 1.8% AMOTA. Further integrating the instance-aware contrastive learning brings an extra 2.1% improvement in AMOTA. Together, our proposed components complement each other to effectively address multi-task learning challenges.
Variants of query filtering. To our knowledge, one-to-many matching has received limited attention in 3D perception. Therefore, we compare our GMM-based filtering module with methods [41, 42, 43] from 2D tasks. Concretely, we divide these methods into IoU-based group and cost-based group. As shown in Tab. 4(b), IoU-based methods fail to surpass even the fixed-threshold method, DETA [43]. We infer that the object queries cannot ensure sufficient IoU overlap with GT boxes. Compared to our dynamic method, DETA proves inadequate for mining potential high-quality candidates. We further demonstrate the candidate selection effect of our GMM-based filtering module in Fig. 6.
Variants of matching. In Tab. 5, we analyze the impact of different matching variants for object queries in hybrid matching module. We consider from the perspectives of decoder type, matching range, and matching strategy, compared to our baseline (highlighted in gray). Based on the experimental results we draw the following two conclusions: 1) The one-to-many assignment is not applicable to the standard decoder due to the NMS effect of the self-attention mechanism (①vs.②). 2) Object queries in the C-decoder cannot benefit from a broader matching range with all ground-truth targets (④vs.⑤). We attribute this to the reduction of positive sample proportion in contrastive learning, which may diminish the effectiveness of synergistic learning. Moreover, matching solely with New-born GTs offers limited improvement due to the reduction in positive supervision (③vs.⑤).
Table 4: Ablation study on components and query filtering methods. The terms "Det" and "Trk" denote the matching for detection and tracking in Task-specific Hybrid Matching module, respectively. "CL" denotes the Instance-aware Contrastive Learning for object and track queries in C-decoder.
(a) Evaluation of component effectiveness.
# Det Trk CL Tracking Detection
[…]
40.8% 1.343 47.7% 37.8%
42.2% 1.327 49.8% 38.9%
40.9% 1.339 48.1% 37.9%
42.6% 1.286 49.9% 39.9%
44.7% 1.262 49.7% 39.6%
(b) Comparison of query filtering methods.
@@
IoU-based
ATSS [41] 39.3% 1.367 49.9% 47.0% 37.3%
SimOTA [42] 42.8% 1.323 52.5% 49.2% 38.8%
Cost-based
DETA [43] 43.1% 1.320 54.2% 49.4% 38.6%
GMM (ours) 44.7% 1.262 56.5% 49.7% 39.6%
Table 5: Comparision results of matching variants for object queries in Task-specific Hybrid Matching module.
# Decoder Type Matching Range Strategy Tracking Detection
@@
S-decoder New-born GTs one-to-one 40.8% 1.343 50.7% 15288 40398 47.7% 37.8%
S-decoder New-born GTs one-to-many 38.8% 1.357 52.4% 15461 39930 46.5% 37.1%
C-decoder New-born GTs one-to-many 41.1% 1.339 52.4% 14265 40231 47.8% 38.3%
@@
Table 6: Ablation of maximum assignment constraint $K$.
@@
Table 7: Ablation of training sample length $T$.
@@
Hyperparameters. We study two types of hyperparameters, the maximum assignment constraint $K$ per GT in Eq. 9 and the training sample length $T$. For $K$, according to Tab. 7, the best performance is achieved when $K=5$. Possibly, a smaller matching number could lack supervision signals, while a larger matching number could introduce noisy and unreliable candidates. For $T$ in Tab. 7, we observe that when the frame number is increased from 2 to 3, SynCL obtains a boost of $+1.6\%$ AMOTA. However, when $T>3$, SynCL cannot improve the model further. We infer that more consecutive frames lead to more optimization targets and temporal gradients, affecting the training stability. We thus use $K=5$ and $T=3$ as the default setting for all our main results and other ablation studies.